# Alignment-Aware Model Extraction Attacks on Large Language Models

### v1

@@ 7.2.3 Fidelity under Different Victim and Local Models
8 Potential Defenses
Query Detection. One approach to effectively prevent the attack of LoRD is by detecting the distribution of query texts. This is because LoRD, similar to current MEA algorithms, makes no improvements to query samples, indicating that it can be detected by analyzing the statistical information of the adversary’s queries, such as the number of queries, distribution of query contents, and so on. However, this defense is usually resource-consuming, as it requires the LLM provider to store all query texts of each user. Besides, the potential for false positives could adversely affect the user experience.
More Powerful Watermarks. While we highlight the watermark resistance of LoRD, watermarking remains one of the most effective solutions to mitigate MEAs. For example, some model-level watermarks, such as backdoor-based watermarking [24, 34], can effectively certify the theft of DNNs. However, model-level watermarking on LLMs remains preliminary. Besides, this technique might not work when the adversary only steals a subset of knowledge in which no backdoor is embedded.
9 Conclusion
In this paper, we have focused on the extraction problem of commercial large language models. We proposed LoRD, a practical and realistic extraction algorithm which is consistent with the alignment procedure of large language models. Our analysis proved that LoRD can reduce the query time significantly and mitigate the certification of current watermarks naturally, surpassing existing MEA algorithms’ capabilities. Extensive experiments on domain-specific stealing demonstrates the superiority of our method.
Appendix A Visualization of Distributions
We also investigate the probability distributions in the generation procedure among different extraction methods. Specifically, we visualize these distributions for four models, the victim model (GPT-3.5-turbo), the initial local model (llama3-8B), and the learned local models with MLE and LoRD. As plotted in Figure 11, each row in the subfigures refers to the distribution when generating the $i$-th token, with each column element indicating the probability predicted for the corresponding token index. We limit the visualization to no more than five token probabilities as currently only GPT-3.5-turbo provides the token prediction probabilities during generation, with a maximum of 5 candidate tokens  [5].
From Figure 11, we can see that both MLE and LoRD successfully redistribute the generation of the initial local model into a distribution similar to the victim model’s, where probabilities, especially Top-1 tokens, have been well inherited in the extraction. This phenomenon supports our analysis in Theorem 1. However, distributions of MLE extracted models are consistently sharper than LoRD’s, which aligns with our analysis in Section 5.2, where we claim that MLE leads local models to overfit to the preferred sentences (i.e., Top-1 tokens), namely PO, and thus to disrupt the original distributions, leveraging unusual low probabilities for other token indexes. The reason why LoRD can be resistant watermarks, i.e., tokens in Top-1, can also be derived from this discovery.
@@ Appendix A Visualization of Distributions
To compare MLE and LoRD accurately, we quantize the entropy of these distributions, and compute the KL divergence ($\mathbb{D}_{\text{KL}}$), and the Spearman Correlation (Spear. Corr.) with respect to the victim and initial local model. As shown in Table 5, while the MLE extracted model exhibits a lower KL divergence (i.e., high distribution similarity) with the victim model than LoRD’s on the training dataset, its KL divergence becomes comparable to LoRD’s on the test set. Meanwhile, its Spearman correlation significantly decreases from 0.78 to 0.27, which shows that MLE cannot effectively imitate prediction behaviors of the victim model when encountering data beyond the training dataset.
Appendix B Limitations and Future Works
MEAs on Multi-modal Models. While this paper delves into MEAs for large language models, it acknowledges the oversight of the multi-modal attribution of current commercial models [48, 15] that integrate various forms of data such as text, images, voice, and so on. The challenge of extending MEA algorithms to accommodate these models, which requires extra considerations on the unified representation of concepts, remains unexplored. Future work could focus on developing MEA methodologies sensitive to multi-modal data nuances.
Capacities beyond LaViSH Settings. We utilize the LaViSH setting to describe the model capacity of adversaries in our threat model (see Section 3). However, sometimes, the adversary might possess comparable or superior training resources to the victims. Though this paper posits that our MEA algorithms and theoretical analysis are still compatible with such conditions, we concede that concrete experimental validation and results beyond LaViSH settings are not presented here.
Lower-level Extractions. This study evaluates MEAs at the performance level, i.e., it measures the extraction effectiveness simply through task performance metrics, or the similarity of learned distributions to the victim model. This setting is justified, as performance metrics are essential for evaluating task-related knowledge and the practical application of LLMs. However, it does not consider the lower-level similarities between the victim and local models. Can we achieve neuron-level alignments in LLM’s MEAs? How does a LaViSH setting hurt LLM’s MEAs? Is it compatible to extract a MoE (Mix-of-the-Expert) [46] victim model with a dense local model? These questions are not addressed in this research.
Appendix C Proofs of Theorems
C.1 The Deduction of Equation 16
@@
$max_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}{R_{\theta_{\phi}}(\mathbf{x},\hat{\mathbf{y}})}}-\beta\mathbb{D}_{KL}[P_{\theta}(\mathbf{y}|\mathbf{x})||P_{\theta_{init}}(\mathbf{y}|\mathbf{x})]$
$\Rightarrow max_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{R_{\theta_{\phi}}(\mathbf{x},\mathbf{y})}}-\beta[\text{log}P_{\theta}(\mathbf{y}|\mathbf{x})-\text{log}P_{\theta_{init}}(\mathbf{y}|\mathbf{x})]$
$\Rightarrow min_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{-\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y})}}+\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}$
$\Rightarrow min_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{-\text{log}(\text{exp}(\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y})))}+\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}}$
$\Rightarrow min_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{\text{exp}(\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y}))\cdot P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}}}.$
@@
$min_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{\text{exp}(\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y}))\cdot P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}}}$
$\Rightarrow min_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{Z(\mathbf{x})\cdot P_{\theta}(\mathbf{y}|\mathbf{x})}{\text{exp}(\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y}))\cdot P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}}}$
@@
$min_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{Z(\mathbf{x})\cdot P_{\theta}(\mathbf{y}|\mathbf{x})}{\text{exp}(\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y}))\cdot P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}}}-\text{log}Z(\mathbf{x})$
$\Rightarrow min_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{P_{\theta*}(\mathbf{y}|\mathbf{x})}}}-\text{log}Z(\mathbf{x}).$
@@
$min_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{P_{\theta*}(\mathbf{y}|\mathbf{x})}}}-\text{log}Z(\mathbf{x})$ (20)
$\Rightarrow min_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\left[\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{P_{\theta*}(\mathbf{y}|\mathbf{x})}}\right]-\text{log}Z(\mathbf{x})}$
$\Rightarrow min_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\mathbb{D}_{KL}[P_{\theta}(\mathbf{y}|\mathbf{x})||P_{\theta*}(\mathbf{y}|\mathbf{x})]-\text{log}Z(\mathbf{x})}.$

### v2

@@ A.2.3 Fidelity under Different Victim and Local Models
We evaluate the fidelity of extracting different victim models using various pre-trained local models. Specifically, we select GPT-3.5, GPT-4, and GPT-4o as victim models, and employ five state-of-the-art open-source models, Phi-3 (3.8B), OPT (6.7B), Qwen-2 (7B), Mistral-V3 (7B), and Llama-3 (8B), as local models, as shown in Figure 9.
Horizontally, while GPT-4 exhibits a consistently lower extracted fidelity compared to the other two victim models, vulnerabilities of the three victim models are generally similar. Vertically, fidelity of different local models can be significantly impacted by their performance. For instance, OPT (6.7B) shows a noticeably lower score compared to the other four models, which indicates that the initial performance of the local model will affect the performance of MEAs. Besides, Phi-3 (3.8B) achieves a comparable fidelity to larger models like Llama-3 (8B), demonstrating that the size of a local model does not influence final fidelity in domain-specific stealing after 2.7 billion, which corroborates the observation in Appendix A.2.2.
A.3 Visualization of Distributions
We also investigate the probability distributions in the generation procedure among different extraction methods. Specifically, we visualize these distributions for four models, the victim model (GPT-3.5-turbo), the initial local model (llama3-8B), and the learned local models with MLE and LoRD. As plotted in Figure 10, each row in the subfigures refers to the distribution when generating the $i$-th token, with each column element indicating the probability predicted for the corresponding token index. We limit the visualization to no more than five token probabilities as currently only GPT-3.5-turbo provides the token prediction probabilities during generation, with a maximum of 5 candidate tokens  (OpenAI, 2024).
From Figure 10, we can see that both MLE and LoRD successfully redistribute the generation of the initial local model into a distribution similar to the victim model’s, where probabilities, especially Top-1 tokens, have been well inherited in the extraction. This phenomenon supports our analysis in Proposition 2. However, distributions of MLE extracted models are consistently sharper than LoRD’s, which aligns with our analysis in Section 4.2, where we claim that MLE leads local models to overfit to the preferred sentences (i.e., Top-1 tokens), namely PO, and thus to disrupt the original distributions, leveraging unusual low probabilities for other token indexes. The reason why LoRD can be resistant watermarks, i.e., tokens in Top-1, can also be derived from this discovery.
     Method BLEU BERTScore Rouge-L
1 2 3 4 Precision Recall F1 F1
LoRD   (Q=16)   (T=0.8) 54.40 42.18 33.56 27.06 89.89 94.06 91.44 56.09
SimPO (Q=16)   (T=1.0) 44.80 34.80 27.94 22.83 89.79 93.50 91.57 48.39
SimPO (Q=16)   (T=1.3) 44.19 33.45 26.31 21.18 88.49 92.65 90.47 47.09
SimPO (Q=16)   (T=0.8) 42.99 31.81 24.85 19.82 90.37 88.32 92.64 44.04
SimPO (Q=256) (T=1.3) 3.09 0.13 0.00 0.00 68.04 81.54 74.17 11.22
SimPO (Q=256) (T=0.8) 20.99 10.75 7.01 5.04 85.56 87.52 86.50 21.08
Table 4: Comparison between LoRD and The Direct Prompting with SimPO. $T$ denotes the temperature of local model’s sampling, and $Q$ denotes the query times.
@@ A.3 Visualization of Distributions
To compare MLE and LoRD accurately, we quantize the entropy of these distributions, and compute the KL divergence ($\mathbb{D}_{\text{KL}}$), and the Spearman Correlation (Spear. Corr.) with respect to the victim and initial local model. As shown in Table 5, while the MLE extracted model exhibits a lower KL divergence (i.e., high distribution similarity) with the victim model than LoRD’s on the training dataset, its KL divergence becomes comparable to LoRD’s on the test set. Meanwhile, its Spearman correlation significantly decreases from 0.78 to 0.27, which shows that MLE cannot effectively imitate victim model’s prediction behaviors when encountering data beyond the training dataset.
A.4 Ablation Study
We conduct an ablation study to assess the impact of our proposed loss functions shown in Section 3. Specifically, we adopt the same experimental settings described in Section 5.1 and compare LoRD against the following variations on the WMT16 (de-en) dataset:
$\bullet$ w.o. $\sigma(\cdot)$: Removing the sigmoid function in Equation 11;
$\bullet$ Rep. $\mathbf{y}^{-}$ w. $\mathbf{y}^{+}$: Replacing $\mathbf{y}^{-}_{t-1}$ with $\mathbf{y}^{+}_{t-1}$ defined in Equation 9;
$\bullet$ w.o. $\mathcal{L}_{reg}$: Eliminating the regularization term.
The ablation results are presented in Table 6. Our findings indicate that the sigmoid function used for normalization is not essential for the effectiveness of our extraction strategy. However, the regularization term proves to be crucial for ensuring the model’s convergence, which is consistent with our theoretical analysis.
A.5 LoRD versus Direct Prompting
We notice that the victim model can serve as a feedback signal to explicitly determine $\mathbf{y}^{+}_{t-1}$ and $\mathbf{y}^{-}_{t-1}$, thereby enabling a reinforcement learning (RL) approach based on direct prompting. This idea aligns with prior work on reinforcement learning with AI feedback (RLAIF), as discussed in Appendix C.1.
In this section, we present an empirical comparison between LoRD and direct prompting and argue that direct prompting is less suitable for MEAs than our LoRD.
     Method BLEU-4 BERTScore Rouge-L
Precision Recall F1 F1
LoRD 27.06 89.89 94.06 91.44 56.09
w.o. $\sigma(\cdot)$ 23.77 89.25 93.73 91.38 50.39
Rep. $\mathbf{y}^{-}_{t-1}$ w. $\mathbf{y}^{+}_{t-1}$ 25.87 87.41 93.28 90.19 54.12
w.o. $\mathcal{L}_{reg}$ NC NC NC NC NC
Table 6: Ablation Study for LoRD. NC denotes that the model does not converged during training.
Empirical Comparison. We design a prompt to obtain feedback from the victim model as “For a translation task involving the conversion of the given ‘Text‘ into English, the user will provide two translation versions labeled ‘A‘ and ‘B‘. Your task is to return the *letter corresponding to the better translation* without including any additional output.”. For direct prompting, we allow the victim model to determine the positive and negative responses generated by the local model. These responses are then used to fine-tune the local model using a DPO-inspired loss function. Specifically, we employ SimPO Meng et al. (2024) as the loss function. To ensure a fair comparison, we maintain the same hyperparameter settings as in previous experiments.
As shown in Table 4, we conducted experiments with various sampling temperatures for the direct prompting. However, the performance of the direct prompting still underperforms LoRD. This limitation may stem from the local model’s lack of guidance from correct answers. When the local model generates two suboptimal responses, a direct prompting-based method is compelled to select the "winner" of two inadequate response rather than an optimal response, which we believe is the crux of the issue.
  BLEU BERTScore Rouge-L
1 2 3 4 Pre. Rec. F1. Pre. Rec. F1.
Text to SQL: WikiSQL (Zhong et al., 2017) with 64 query samples
Victim Model 54.1 41.4 32.1 24.4 86.9 93.5 90.1 58.9 62.1 59.7
Text to SQL: Spider (Zhong et al., 2017) with 64 query samples
Victim Model 9.4 3.9 2.1 1.1 77.7 84.1 80.6 17.1 36.3 21.8
Data to Text: E2E NLG (Dušek et al., 2020) with 64 query samples
Victim Model 51.8 27.0 26.8 19.1 93.9 94.6 94.2 49.6 54.6 51.4
Data to Text: CommonGen (Lin et al., 2020) with 64 query samples
Victim Model 33.3 18.5 11.1 6.9 91.3 92.1 91.7 33.6 40.7 36.1
Summarization: TLDR (Kirk et al., 2023) with 64 query samples
Victim Model 11.9 5.0 2.6 1.5 85.9 88.4 87.1 13.4 30.9 18.4
Summarization: CNN Daily Mail (Hermann et al., 2015) with 64 query samples
Victim Model 20.4 10.8 6.4 4.1 86.4 87.8 87.1 22.4 40.8 28.2
Summarization: Samsum (Gliwa et al., 2019) with 64 query samples
Victim Model 20.7 11.4 6.9 4.4 88.1 91.7 89.8 24.2 50.5 31.6
Table 7: MEA comparison on three tasks, including structured text generation, data to text, and summarization. We use GPT-3.5-turbo as the victim model, and Llama3-8B (Grattafiori et al., 2024) as the local initial model. The intensity of the red or blue color corresponds to the degree of underperformance or outperformance relative to the victim model.
RLHF tackles this challenge by incorporating a regularization term with the initial model, LoRD addresses it through our $\mathcal{L}_{reg}$, and DPO resolves it by employing the training corpus of the reward model. Unfortunately, a direct prompt-based method overlooks this point. To further investigate this problem, we increased the query number to 256, which resulted in the local model failing to converge and exhibiting poor performance.
Besides, we also observed a bias in the victim model’s selection between the first and second sentences. In a series of 256 queries, the model successfully provided an answer (either A or B) 255 times. However, it chose the first sentence only 84 times, which is a mere 32.94%, significantly deviating from the expected 50%. Given that the generated sentences are randomly sampled from the local model without any significant correlation to their order, we deduce that relying on the victim model to directly generate feedback might be, at least, an unreliable approach. It may necessitate additional considerations for the design of the prompt and the capabilities of the victim model to ensure the robustness of these algorithms.
Discussion on the Feasibility. In addition to the empirical comparison, we provide a discussion supporting the proposition that direct prompting is unsuitable for model extraction attacks for the following reasons:
  • A direct feedback query will expose the intention of the adversary;
  • Unlike the current design of LoRD, direct feedback is contingent upon the local model’s responses, which is query-inefficient. Specifically, for a given query sample, the algorithm would need to repeatedly query the victim model to distinguish between $\mathbf{y}^{+}_{t-1}$ and $\mathbf{y}^{-}_{t-1}$ across different learning periods. On the contrary, LoRD necessitates only a single query per sample to discriminate different $(\mathbf{y}^{+}_{t-1},\mathbf{y}^{-}_{t-1})$ pairs;
  • The threat model will change if employing a direct prompting. As we know, both LoRD and MLE are currently trained under the same conditions, i.e. $(\mathbf{x},\mathbf{y}_{vic})$ paires. The fairness would be questioned when we compare methods under disparate query settings.
Appendix B Theoretical Explanations and Proofs
B.1 Explanation of Proposition 1
[…]
Expected Distribution of KD. Following a previous work (Hinton et al., 2015), the objective function of KD is
where $\text{SM}(\cdot)$ represents the softmax function, and $T>1$ denotes the temperature to smooth the targeted distribution $P_{\theta_{vic}}(\cdot|\mathbf{x})$. As described in Equation 14, knowledge distillation aims to align $P_{\theta}(\cdot|\mathbf{x})$ with $P_{\theta_{vic}}(\cdot|\mathbf{x})$ in both the original and the smoothed probability across all dimensions, which is exceptionally comprehensive among these MEA loss functions.
where $\theta*$ denotes the expected parameters of the models as
We provide a detailed derivation for Equation 16 in Appendix B.2. By replacing Equation 15 with Equation 16, the expected distribution can be represented as $\mathbf{r}_{i,j}\cdot P_{\theta_{init}}(\cdot|\mathbf{x})$, in which $\mathbf{r}_{i,j}$ indicates the wrapped distribution gain. This distortion aims to maximize the ratio $P_{\theta}(y^{+}_{j}|\mathbf{x},\mathbf{y}^{+}_{<j})/P_{\theta}(y^{-}_{j}|\mathbf{x},\mathbf{y}^{-}_{<j})$, and leave the probabilities in other dimensions unconstrained directly.
B.2 The Deduction of Equation 16 in Proposition 1
@@
$\max\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}{R_{\theta_{\phi}}(\mathbf{x},\hat{\mathbf{y}})}}-\beta\mathbb{D}_{KL}[P_{\theta}(\mathbf{y}|\mathbf{x})||P_{\theta_{init}}(\mathbf{y}|\mathbf{x})]$
$\Rightarrow\max\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{R_{\theta_{\phi}}(\mathbf{x},\mathbf{y})}}$
$\quad\quad\quad-\beta[\text{log}P_{\theta}(\mathbf{y}|\mathbf{x})-\text{log}P_{\theta_{init}}(\mathbf{y}|\mathbf{x})]$
$\Rightarrow\min\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{-\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y})}}+\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}$
$\Rightarrow\min\limits_{\theta}\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{-\text{log}(\text{exp}(\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y})))}$
$\quad\quad\quad+\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}$
$\Rightarrow\min\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{\text{exp}(\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y}))\cdot P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}}}.$
@@
$\min\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{\text{exp}(\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y}))\cdot P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}}}$
$\Rightarrow\min\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{Z(\mathbf{x})\cdot P_{\theta}(\mathbf{y}|\mathbf{x})}{\text{exp}(\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y}))\cdot P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}}}$
@@
$\min\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{Z(\mathbf{x})\cdot P_{\theta}(\mathbf{y}|\mathbf{x})}{\text{exp}(\frac{1}{\beta}R_{\theta_{\phi}}(\mathbf{x},\mathbf{y}))\cdot P_{\theta_{init}}(\mathbf{y}|\mathbf{x})}}}$
$\quad\quad\quad-\text{log}Z(\mathbf{x})$
$\Rightarrow\min\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{P_{\theta*}(\mathbf{y}|\mathbf{x})}}}-\text{log}Z(\mathbf{x}).$
@@
$\min\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{P_{\theta*}(\mathbf{y}|\mathbf{x})}}}-\text{log}Z(\mathbf{x})$ (18)
$\Rightarrow\min\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\left[\sum_{\mathbf{y}\sim P_{\theta}(\cdot|\mathbf{x})}{\text{log}\frac{P_{\theta}(\mathbf{y}|\mathbf{x})}{P_{\theta*}(\mathbf{y}|\mathbf{x})}}\right]-\text{log}Z(\mathbf{x})}$
$\Rightarrow\min\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}\mathbb{D}_{KL}[P_{\theta}(\mathbf{y}|\mathbf{x})||P_{\theta*}(\mathbf{y}|\mathbf{x})]-\text{log}Z(\mathbf{x})}.$